# Visual Test-time Scaling for GUI Agent Grounding

### v1

@@ Results
Figure 7 shows the distribution of step differences between the combined BaseModel + RegionFocus approach and the BaseModel alone over 400 trajectories. Only the actual browser-interactive steps are counted, excluding RegionFocus overhead. Here the BaseModel is UI-TARS-72B. As shown, BaseModel + RegionFocus generally yields more steps on average ($19.74$% steps), correlating with a overall $34.3$% higher success rate. RegionFocus is triggered $5.8$ times on average for each trajectory in Web Browsing. On average, RegionFocus is triggered $5.8$ times per Web Browsing trajectory. Furthermore, in $32.3$% of cases RegionFocus is triggered only once, yet a single trigger yields an impressive $83.7$% increase in the success of those trajectories.
@@ Appendix A More experimental details
Prompt for Action Prediction – UI-TARS

### v2

@@ Results
Figure 7 shows the distribution of step differences between the combined BaseModel + RegionFocus approach and the BaseModel alone over 400 trajectories. Only the actual browser-interactive steps are counted, excluding RegionFocus overhead. Here the BaseModel is UI-TARS-72B. As shown, BaseModel + RegionFocus generally yields more steps on average ($19.74$% steps), correlating with a overall $34.3$% higher success rate. On average, RegionFocus is triggered $5.8$ times per Web Browsing trajectory. Furthermore, in $32.3$% of cases RegionFocus is triggered only once, yet a single trigger yields an impressive $83.7$% increase in the success of those trajectories.
@@ 4.3 Ablation Studies
4.4 More Analysis
Runtime Analysis
For each trigger, the VLM performs one inference for focal-point proposal, four inferences for region action prediction (which can be executed in parallel), and one inference for action aggregation. The four region action prediction inferences result from adopting four predefined bounding boxes around each predicted focal point, a hyper-parameter. For the interactive environment (WebVoyager), RegionFocus occurs in $61.7\%$ of all trajectories; among these, it is triggered an average of $5.84$ times, incurring an average overhead of $66.8\%$ per trajectory. Notably, $32.3\%$ of triggered cases occur only once, with this single trigger yielding an $83.7\%$ improvement in success. For the static environment (ScreenSpot-Pro), RegionFocus occurs in $60.2\%$ and $33\%$ of cases for the 72B and 7B models, respectively. The 72B model incurs overheads of $180\%$ when region predictions are executed in parallel, and $360\%$ if executed sequentially. Regarding memory usage, RegionFocus requires no additional GPU memory if region predictions are performed sequentially without employing SAM [16]. Whether region action predictions should be executed sequentially or in parallel, the number of regions to zoom into, and whether to utilize SAM depend on the specific use case.
Failure Case Analysis
According to our observations, we summarize the main causes of failures as follows: (1) Element not visible: In cases where the correct element is not visible on the current page (e.g., requires scrolling actions), zooming in on the current page is not beneficial. (2) Non-interactable elements: Some webpages contain extensive text that appears relevant but is not clickable, causing RegionFocus to zoom into these regions and subsequently fail to proceed due to the lack of interactable elements. For instance, on the Hugging Face homepage (https://huggingface.co/), there is abundant relevant text such as “Image-to-Text” and “Sentence Similarity,” which relates to agent tasks but is non-clickable. (3) Wrong focal points: In some rare cases, RegionFocus proposes the wrong focal point, directing attention to irrelevant regions. (4) Action prediction failure: Occasionally, despite identifying the correct focal point, all regional action predictions fail; our observations suggest that employing SAM [14] could alleviate this issue as it further reduces the background clutter. (5) Action aggregation failure: Among the regional actions, the selected one may not be optimal. (6) Others: Other errors arise from general reasoning failures, bot detection, or reaching the maximum step limit.
@@ Appendix A More experimental details
Prompt for Region Focus